Two identical sphere carry q and -2q charge at distance d , when touched and brought back at same position find new repulsive force?

Grade:12th pass

1 Answers

Sagnik jana
6 years ago
Initially,Net charge= q +(-2q)=-qFinally,Charge on first sphere=-q/2Charge on second sphere=-q/2Force=-kq^2/4d^2